when fourteen is reduced by two-thirds of a number, the result is 4. find the number
Answer by macston(5194)

14-(2/3)n=4 Add (2/3)n to each side.
14=2/3n+4 Subtract 4 from each side.
10=2/3n Multiply each side by 3/2
30/2=n
15=n ANSWER: The number is 15.